# OptiChat: Bridging Optimization Models and Practitioners with Large Language Models

### v1

@@ 4.2 Optimization Fundamentals
where $\mathbf{\delta A^{+}}$, $\mathbf{\delta A^{-}}$, $\mathbf{\delta b^{+}}$ are nonnegative slack variables. Since all the constraints are inequalities, adding nonnegative slacks $\mathbf{\delta b^{+}}$ on the right-hand side relaxes the problem. In contrast, adding the left-hand side slacks, $\mathbf{\delta A^{+}}$ and $\mathbf{\delta A^{-}}$, are less desirable. It is worth noting that because only a subset of parameters can be adjusted in practice, the slack variables differ in dimensionality from the parameters $\mathbf{A}$ and $\mathbf{b}$. The objective is to minimize the total perturbation to the original problem by summing up all the slack variables. In principle, different weights could be assigned to different slack variables, representing the cost of perturbing the corresponding parameters.
Sensitivity Analysis OptiChat performs sensitivity analysis based on well-established duality theory in linear programming (LP) [48] where $\mathbf{x}$ does not contain integer variables. In short, the change in the optimal objective value in response to changes in input parameters can be expressed as a value function
@@
where $\mathcal{X}=\{\mathbf{x}:\mathbf{A}\mathbf{x}\leq\mathbf{b}\}$ and $\mathbf{y^{*}}^{\top}$ is the optimal solution to the dual problem of (1). Problem (1) called a primal problem that finds the minimum cost while satisfying all the constraints, whereas the dual problem of (1) aims to seek the tightest lower bound on the optimal primal cost. Denote the original data given in the primal problem as $\mathbf{b^{*}}$, the expression (3) indicates that when the perturbed parameters $\mathbf{b}$ are close to $\mathbf{b^{*}}$, the optimal primal objective value is a linear function of the input parameters with gradient $\mathbf{y^{*}}$. This suggests that the optimal dual solution for the problem with original input parameters can be used to infer the sensitivity of the optimal objective value in the vicinity of $\mathbf{b^{*}}$. The access to dual information in LP is supported by most optimization solvers [41, 46, 49].
Model Modification For what-if queries, the model (1) can be further revised by updating parameters and constraints:
@@
Counterfactual Explanation A counterfactual explanation addresses “why-not” queries by examining alternative decisions or outcomes that are not currently realized in the optimal solution. For example, one might ask: “Why not choose supplier 1?” in a supply chain model, even though the current optimal solution does not include it. To investigate such a counterfactual scenario, we modify the original model by incorporating an additional constraint that forces the desired alternative to occur. In this example, if $x_{1}$ represents the binary decision to select supplier 1, we add the constraint $x_{1}=1$ to enforce this choice. Once we re-solve the modified problem with this counterfactual constraint in place, we can observe how the objective value and the feasibility of the solution change. For instance, the forced selection of supplier 1 may raise production costs, reduce overall profitability, or even make the problem infeasible. By comparing the new solution to the original one, we gain insight into the trade-offs and underlying reasons that the original optimal solution did not include supplier 1. This counterfactual reasoning thus provides a transparent and actionable explanation of the model’s decision-making process, illustrating which model parameters and constraints are most influential in ruling out the queried alternative.
@@ 4.3 Exceptions Management
When practitioners interact with OptiChat, the system is robust in handling exceptions caused by the lack of optimization expertise. In diagnosing queries, adding slack variables to the left-hand side parameters $\mathbf{A}$ will lead to a product of variables between $\mathbf{\delta A^{+}}$, $\mathbf{\delta A^{-}}$ and $\mathbf{x}$ in (2). This results in a non-convex mixed-integer quadratically constrained program (MIQCP), which is often prohibitive to solve. In many situations, left-hand side parameters represent immutable properties. For example, in the constraint $\mathbf{A}\mathbf{x}\leq\mathbf{b}$ where $\mathbf{b}$ is a deadline, $\mathbf{x}$ indicates task status, and $\mathbf{A}$ represents processing times. The processing times are an inherent property of the machines and cannot be changed. When OptiChat detects a request to add slacks to $\mathbf{A}$, it will alert users of this immutability before initiating the MIQCP.

### v2

@@ 3.1.2  Diagnosing Query
where $\mathbf{\delta A^{+}}$, $\mathbf{\delta A^{-}}$, $\mathbf{\delta b^{+}}$ are nonnegative slack variables. Since all the constraints are inequalities, adding nonnegative slacks $\mathbf{\delta b^{+}}$ on the right-hand-side relaxes the problem. In contrast, adding the left-hand-side slacks, $\mathbf{\delta A^{+}}$ and $\mathbf{\delta A^{-}}$, are less desirable. It is worth noting that because only a subset of parameters can be adjusted in practice, the slack variables differ in dimensionality from the parameters $\mathbf{A}$ and $\mathbf{b}$. The objective is to minimize the total perturbation to the original problem by summing up all the slack variables. In principle, different weights could be assigned to different slack variables, representing the cost of perturbing the corresponding parameters. The optimal solution found in (2) can be explained as the minimal adjustments needed to restore feasibility.
It should be noted that adding slack variables to the left-hand-side parameters $\mathbf{A}$ will lead to a product of variables between $\mathbf{\delta A^{+}}$, $\mathbf{\delta A^{-}}$ and $\mathbf{x}$ in (2). This results in a non-convex mixed-integer quadratically constrained program (MIQCP), which is often prohibitive to solve. In many situations, left-hand-side parameters represent immutable properties. For example, in the constraint $\mathbf{A}\mathbf{x}\leq\mathbf{b}$ where $\mathbf{b}$ is a deadline, $\mathbf{x}$ indicates task status, and $\mathbf{A}$ represents processing times. The processing times are an inherent property of the machines and cannot be changed. When OptiChat detects a request to add slacks to $\mathbf{A}$, it will alert users of this immutability before initiating the MIQCP.
3.1.3  Retrieval Query
An optimization model consists of decision variables, parameters data, constraints, and an objective function. Optimization models developed for practical use are often large in scale, with some components indexed over large sets. For example, a variable that represents the assignment decision of an aircraft can be indexed over hundreds of routes. When practitioners need to review specific data or optimal decisions from the model, it is inefficient to retrieve this information. OptiChat facilitates real-time access to specific model information through natural language.
3.1.4  Sensitivity Query
Many optimization models are constructed with incomplete knowledge of problem parameters. For example, electricity prices and customer demands often fluctuate in the market over time and cannot be fixed within the model. The values of these parameters can be predicted using historical data and updated on a rolling basis. It is important for practitioners to evaluate how changes in problem parameters impact the optimal objective value, undertake risk assessment, and devise appropriate management strategies.
We propose to perform sensitivity analysis based on well-established duality theory in linear programming (LP) [55] where $\mathbf{x}$ does not contain integer variables. In short, the change in the optimal objective value in response to changes in input parameters can be expressed as a value function
@@ 3.1.4  Sensitivity Query
where $\mathcal{X}=\{\mathbf{x}:\mathbf{A}\mathbf{x}\leq\mathbf{b}\}$ and $\mathbf{y^{*}}^{\top}$ is the optimal solution to the dual problem of (1). Problem (1) called a primal problem that finds the minimum cost while satisfying all the constraints, whereas the dual problem of (1) aims to seek the tightest lower bound on the optimal primal cost. Denote the original data given in the primal problem as $\mathbf{b^{*}}$, the expression (3) indicates that when the perturbed parameters $\mathbf{b}$ are close to $\mathbf{b^{*}}$, the optimal primal objective value is a linear function of the input parameters with gradient $\mathbf{y^{*}}$. In other words, it precisely reflects the sensitivity of the optimal objective value in the vicinity of $\mathbf{b^{*}}$. The access to dual information in LP is supported by most optimization solvers [49, 53, 56].
3.1.5  What-if Query
Although sensitivity analysis can evaluate local changes in parameters, it is not a valid methodology to evaluate larger changes. This limitation arises because the optimal dual solution in (3) is not constant but depends on parameters. After large parameter perturbations, the dual solution previously found in the original problem no longer accurately reflects the change in optimal objective value. These significant perturbations often occur when a new policy is being established, or the industry is evaluating a new business strategy. For example, practitioners may pose questions such as “What if we increase the labor force to 35 people?” and “What if customer orders are cut by a third?”
In these cases, OptiChat systematically identifies the extent and type of modification from the user’s query. The model (1) is then revised by updating parameters and constraints accordingly:
@@ 3.1.5  What-if Query
3.1.6  Why-not Query
Mathematical optimization is a rigorous methodology that searches for the optimal solution. In contrast, practitioners often rely on their experience to make decisions. As a result, it can be challenging to convince business managers or stakeholders to accept the optimal solutions given by an optimization model, especially when these solutions are against their intuitions. For example, one might ask: “Why not choose supplier 1?” in a supply chain model, even though the current optimal solution does not include it.
@@ 3.1.6  Why-not Query
A counterfactual explanation addresses the “why-not” queries by examining alternative decisions or outcomes that are not currently realized in the optimal solution. To investigate such a counterfactual scenario, we modify the original model by incorporating an additional set of constraints that force the desired alternative to occur. In this example, if $x_{1}$ represents the binary decision to select supplier 1, we add the constraint $x_{1}=1$ to enforce this choice. Once we re-solve the modified problem with this counterfactual constraint in place, we can observe how the objective value and the feasibility of the solution change. For instance, the forced selection of supplier 1 may raise production costs, reduce overall profitability, or even make the problem infeasible. By comparing the new solution to the original one, we gain insight into the trade-offs and underlying reasons that the original optimal solution did not include supplier 1. This counterfactual reasoning thus provides a transparent and actionable explanation of the model’s decision-making process, illustrating which model parameters and constraints are most influential in ruling out the queried alternative. The explanation for the “why-not” query relies on code generated by the LLM since the counterfactual can be any arbitrary fact suggested by the practitioners, for which predefined functions do not suffice.
3.2 Multi-agent Framework
Figure 3 depicts the workflow of the multi-agent framework. Before the interactive dialogue, OptiChat begins with the Illustrator agent, which preprocesses the Pyomo optimization model uploaded by users. During the conversation, the Coordinator routes queries either to the Explainer agent, when they can be directly answered by an LLM, or to the team of Engineering agents when more specialized reasoning is required. The Engineering team operates in a structured order, comprising the following agents: (1) the Operator, who controls all predefined functions; (2) the Programmer and Evaluator agents, responsible for generating code for tasks that cannot be addressed by the predefined functions; and (3) the Reminder agent, who provides supplementary information to enhance the accuracy of the other two. The prompts are shown in Appendix C. To illustrate the workflow, a what-if query is provided as an illustrative example in Appendix D.
3.2.1  Illustrator
This agent extracts sets, parameters, variables, constraints, and objectives from the Pyomo optimization model, labeling each component with a natural language description. These descriptions establish a dictionary that maps the physical meanings of model components to their corresponding notations in the source code. Since users may refer to model components using different terms, this preprocessing enables LLMs to accurately identify the referenced component by providing the problem context. Alongside these descriptions, information about components, such as index dimensions and solution status, is also stored to support interaction with optimization tools. Finally, the Illustrator agent describes the model to users in a concise and coherent manner using natural language. If the solution status of a model indicates infeasibility, the agent will also invoke tools to compute the IIS and provide troubleshooting recommendations by interpreting the conflicting constraints.
3.2.2  Coordinator
After the Illustrator agent preprocesses a model, users can submit queries to initiate an interactive conversation. The Coordinator agent classifies each query as solution-agnostic or solution-specific. Solution-agnostic queries are solely addressed by the Explainer agent through in-context reasoning, while solution-specific queries require technical feedback from agents in the Engineering team before being forwarded to the Explainer agent.
3.2.3  Explainer
The Explainer agent is responsible for conveying any technical information to practitioners in an understandable manner as the endpoint for all queries.
3.2.4  Reminder
The Reminder agent serves as the entry point for every solution-specific query, guiding LLMs to more accurately determine which function to invoke, which model component to modify, and which specific index within that component to reference. Among these tasks, selecting the correct component index (e.g., a parameter index) is the most error-prone. In contrast, identifying the appropriate function name is generally easier for LLMs, aided by few-shot demonstrations. The model description provided by the Illustrator agent helps the Reminder to pinpoint the relevant component name. The syntax guidance offered by the Reminder agent has been shown in our ablation studies (Section 4.2.4) to improve the accuracy in identifying the correct function, component names and indices.
3.2.5  Operator
As proposed in Section 3.1, we implement four predefined functions in the Operator agent to address the diagnosing, retrieval, sensitivity, and what-if queries in Figure 2. Informed by the specific syntax guidance, this agent selects the appropriate function and arguments, then invokes the corresponding tool to generate the solution.
3.2.6  Programmer And Evaluator
The development of the Programmer and Evaluator agents is inspired by recent works on applying code generation to optimization [16, 18]. The code generation capability of LLMs has been widely explored and demonstrated impressive results across various tasks [57]. These agents are designed to address the why-not queries in Figure 2, but also potentially handle unexpected queries beyond the reach of predefined functions. The Programmer and Evaluator are involved in a loop with limited iterations. The Evaluator first executes the code generated by the Programmer, then reviews the terminal outputs and error messages if available. Throughout this process, the code is automatically refined until a bug-free and comprehensive solution is produced. The prompts for why-not queries are designed to guide the LLM in generating additional constraints from the user’s counterfactual query, thereby narrowing the LLM’s task scope and reducing its code output.